Grade 11Mechanics

The gravitational potential energy of a body at a distance r from the centre of earth is U. Its weight at a distance 2r from the centre of earth is

a) U/r

b) U/2r

c) U/4r

d) U/root[2]r

Arnab Mandal
14 Years agoGrade 11

2 Answers

Ashwin Muralidharan IIT Madras
ApprovedApproved Tutor Answer14 Years ago

Hi Arnab,

 

The Gvt Potential Energy at a distance R is,

 

U = GMm/r

 

And weight at distance 2r is

 

W = GMm/(2r)2 = GMm/4r2 = U/4r.

 

Hence Option (C).

 

Hope this helps.

 

Best Regards,

Ashwin (IIT Madras).

nikhil manali
14 Years ago

no since u=GMm/r

and GM/r2=g                                                 r2 is r square

so GM=gr2

therefore  u=gr2m/r

and mg= w

u=wr                since r2 and r wiil be divedid by each other

now r=2r

so w=u/2r